OpenStudy (anonymous):

cube root of 7 multiplied by the square root of 7 over sixth root of 7 to the power of 5 Solve.. Please help!!!

OpenStudy (anonymous):

\[^3\sqrt7\sqrt7/^6\sqrt7^5\]

hartnn (hartnn):

lets first use this : \(\huge \sqrt[n]x = x^{1/n}\) so what about \(\large \sqrt[3]7 = ... ?\)

OpenStudy (anonymous):

would it be \[7^1/3\]

OpenStudy (anonymous):

1/3 is supposed to be in exponent form

hartnn (hartnn):

yes, i know :) then you can also answer me \(\sqrt[6]7^5 = ... \\ \sqrt 7 = ...\)

OpenStudy (anonymous):

\[^6\sqrt7^5 = 7^5/6\] and \[\sqrt7\] stays the same right? I am sorry, but I am not really good at this....

hartnn (hartnn):

you're good! both of your answers are correct :) and for \(\sqrt 7\) it becomes \(\Large 7^{1/2}\)

OpenStudy (anonymous):

oh ya....

hartnn (hartnn):

now moving to next rule : \(\huge x^m x^n = x^{m+n}\) so